### v1

# The Role of Inclusion, Control, and Ownership in Workplace AI-Mediated Communication

Abstract.
Large language models (LLMs) can exhibit social biases. Given LLMs’ increasing integration into workplace software, these biases may impact workers’ well-being and may disproportionately impact minoritized communities. This short paper investigates how co-writing with an LLM impacts three measures related to user’s well-being: feelings of inclusion, control, and ownership over their work. In an online experiment, participants wrote hypothetical job promotion requests to their boss and using either hesitant or self-assured auto-complete suggestions from an LLM. Afterward, participants reported their feelings of inclusion, control, and ownership. We found that the style of the AI model did not impact perceived inclusion. Furthermore, individuals with higher perceived inclusion also perceived greater agency and ownership, an effect more strongly impacting participants of minoritized genders. Lastly, feelings of inclusion can mitigate a loss of control and agency when accepting more AI suggestions. Future work should explore feelings of inclusion in AI-written communication.
1. Introduction
Large language models (LLMs) have the potential to transform workplace communication. These models are already seeing widespread use assisting in common workplace tasks like writing emails. Recent scholarship has shown that workers can use LLMs at work to increase their productivity and produce higher-quality work (Noy and Zhang, 2023). However, as others have noted, one potential challenge is that language models can reflect social biases in the language they produce (Weidinger et al., 2021; Bender et al., 2021). While the existence of such biases is well-documented, we know far less about how these biases affect individual users. In particular, when AI tools—which may reflect prevailing social norms—are used in the workplace, they may inflict a “psychological tax” for individuals who do not align with these norms (Weidinger et al., 2021), affecting their perceived inclusion, control, and ownership over their work output.
Feelings of inclusion, control, and ownership are core to the human experience (Ryan and Deci, 2017). People who feel included develop a sense of belonging (Sargent et al., 2002) that can significantly improve their psychological health (Lambert et al., 2013; Reker and Wong, 2013). For example, people with a strong sense of belonging have shown increased resilience (Sargent et al., 2002), and a strong sense of belonging in the workplace can mediate negative feelings like anxiety and depression (Hartley, 2011; Cockshaw and Shochet, 2010). Furthermore, employees with a strong sense of belonging are more motivated to be productive and reach their fullest potential (Isham et al., 2020; Walton et al., 2012; Waller, 2020). Similarly, perceptions of control and ownership positively contribute to employees’ workplace satisfaction and job performance (Dyne and Pierce, 2004; Eatough and Spector, 2014; Pierce et al., 2004).
Given the relationship between inclusion, control, and ownership and their implications for individual well-being in professional environments, it is crucial to explore how elements of workplace AI-mediated communication (AI-MC) (Hancock et al., 2020) can impact these feelings, especially for users’ from minoritized groups. In particular, we focus here on the potential impact of AI writing assistants that provide auto-complete suggestions as the user is typing text, for example while writing an email. In this context, we ask the following questions:
RQ1: Can the style of AI auto-complete suggestions impact people’s sense of inclusion, control, and ownership over the text they write?
RQ2: How do these perceptions differ by gender?
In the current study, we investigate the impact of a specific communication style— assertiveness—on minoritized gender groups. We build on a rich history examining gender disparities, particularly in the workplace where assertive language can be very consequential (Amanatullah and Morris, 2010; Smith, 2002) and associated with better work outcomes (Mazei et al., 2015; Amanatullah and Morris, 2010). While there are many cultural differences that guide expectations of gender presentation in the workplace, scholars in gender-based communication describe male talk in Western workplaces as being assertive, powerful, and authoritative (Mullany, [n. d.]; Merchant, 2012; Thimm et al., 2008). If writing assistants making suggestions in a more assertive style are not as well-aligned with the communication style of women and members of other minoritized gender groups, the discrepancy in style alignment could cause those individuals to feel a lower sense of inclusion, control, and ownership.
Specifically, this short paper addresses our research questions using an online experiment where participants completed a hypothetical, but common and consequential workplace writing task, namely writing to their manager asking for a promotion. Participants received auto-complete suggestions from one of two LLM-powered AI writing assistants that offered suggestions in different styles: assertive or hesitant. We then asked participants about their feelings of inclusion, control, and ownership in the writing experience.
We found that the intervention—getting suggestions in assertive or hesitant style from an AI assistant—did not impact participants’ perceptions of inclusion, though it did impact their perceptions of control and ownership. Surprisingly, we did not find the assistant style treatments directly resulted in differences between genders in our main measures. We do find evidence that gender plays some role: namely that minoritized genders’ perception of control and ownership are more greatly impacted by their (likely pre-existing) feelings of inclusion than men’s. In light of these findings, we offer a new conceptual model for understanding the impact of these factors on people’s feelings of overall agency in AI-mediated environments.
2. Background
While AI-powered intelligent writing assistants are a relatively recent development, there is already a rich set of research contributions on our interactions with writing assistants and the biases they may introduce.
As large language models became more robust, with the ability to inspire ideas, revise text, and generate short stories, the field has shifted toward viewing other dimensions of the writing process (Singh et al., 2022; Bhat et al., 2023). There is a growing interest in viewing writing assistants powered by large language models as co-authors and studying the interaction between users and these systems (Lee et al., 2022; Mirowski et al., 2023). Recent scholarship in this area has focused on understanding how writers evaluate and integrate the suggestions provided by these language models into their cognitive writing processes (Bhat et al., 2023). Writers proactively engage with system-generated content, making deliberate efforts to integrate it into their writing (Singh et al., 2022). However, there is evidence that perceptions of ownership can differ. Inexperienced writers perceive less ownership than experienced writers during the co-writing process (Mieczkowski, 2022). Here, we expand the understanding of how these systems can best support and enhance the writing process by demonstrating the importance of inclusion when interacting with an AI writing assistant.
As large language models like GPT (Vaswani et al., 2023; Bommasani et al., 2022) are increasingly integrated into our society, there is growing concern among AI ethicists regarding their potential harms, including in the context of auto-complete suggestions. Concerns have been raised regarding the risks of misinformation, privacy violations, socioeconomic harms, and, most salient to our work here, representational harms (Weidinger et al., 2021). Large language models are trained on large corpuses of data which often reflect historic and systemic injustices. For example, a language model is more likely to produce the phrase “he is a doctor” than “she is a doctor”, assuming the subject’s gender or ethnic identity (Weidinger et al., 2021). As a result, these models may produce demeaning language and stereotypes which can be especially harmful for some with intersecting identities. Specifically in our context, biases in auto-complete suggestions have been shown to shift the use of language and influence content (Hohenstein et al., 2023; Poddar et al., 2023; Jakesch et al., 2023). AI suggestions even shift people’s opinions (Jakesch et al., 2023). Extending this body of work, here we show how AI biases may impact the perceptions of people using these technologies.
3. Methods
We addressed our research questions using an online experiment where participants completed a common workplace writing task. We asked participants to write a hypothetical email to their manager asking for a promotion. We chose the promotion task since it is a situation that requires high assertiveness, a context in which minoritized genders face several challenges (Amanatullah and Morris, 2010; Amanatullah and Tinsley, 2013; Mazei et al., 2015). The experiment had three conditions. Participants in two of the treatments received auto-complete suggestions from a language model. In one of these treatments, the model made assertive auto-complete suggestions, and in the other, it made hesitant auto-complete suggestions. A control group of participants wrote without the assistance of any model.
After the writing task, participants answered a demographic questionnaire. Participants in the AI conditions answered an additional post-task questionnaire regarding their perceptions of inclusion, agency, and ownership while writing with the AI assistant. The questions and 5-point Likert scale answers were adapted from previous work (Mieczkowski, 2022; Jakesch et al., 2023). The research design was approved by [anonymized] IRB. The design was preregistered and is available (anonymously) on OSF111https://osf.io/kwbev?view_only=c5639dcc6292482b8bd79f2086b38d24.
3.1. Writing App
We used a custom experimental platform combining a rich-text editor and a writing assistant, as seen in Figure 1. The platform code was made available by researchers (Jakesch et al., 2023). The suggestions work such that when the participant pauses typing, the system displays suggestions that are 20 words or less. The user can then choose to accept the suggestion by pressing tab or right arrow key, or they can ignore the suggestion by continuing to type. Pressing tab accepts the subsequent word in the suggestion, and a user can press multiple times to accept the full suggestion.
To generate the text suggestions for the experiment in real time, we used GPT-4 via an API call with prompts carefully designed to display assertive or hesitant text. To achieve that goal, we inserted the prompt “In a [self-assured/hesitant] manner ask for a raise” before the participants’ written text. After the entered text, we added the affirming instruction “add one sentence continuing the email in a [self-assured/hesitant] manner asking for a raise”. This way, the suggestions maintained the desired style regardless of the text already entered by the participant.
3.2. Participant Recruitment
We recruited 750 participants from a gender-balanced sample of English-speaking US-based adults on Prolific. The sample size was calculated using a two-way ANOVA with a power of 80% based on a pre-test survey. Tp ensure response quality, we required participants to have approval ratings greater than 95% on Prolific. We excluded participants who did not state their gender, and manually verified that each writing sample was faithful to the writing task. In the end, we had usable data from 738 participants.
Our experiment placed emphasis on two distinct gender categories: individuals with significant societal influence, predominantly men, and minoritized genders who typically encounter societal disparities. While describing our results we will refer to Women, Non-Binary or third gender, and “preferred to describe” (also known as Self-Described) as WNBSD participants. Of the 738 participants whose data is used, 48.9% identified as men, and 51.1% as WNBSD: 47.7% identified as women, 3% identified as non-binary or third gender, and 0.4% preferred to describe their gender.
The majority of the participants were between 25-34 years old (56%), identified as white (61.9%), received a bachelor’s degree (38.2%), reported working a full-time job (59.4%), and reported working entirely in person (40.6%).
3.3. Measures
We used the following measures to examine how participants interacted with the AI assistant and their perceptions of the writing task with the AI assistant.
AI reliance. Prior work on how users interact with AI writing assistants uses the measure of “mutuality” to describe the level of interaction the writer has with the AI assistant (Mieczkowski, 2022; Lee et al., 2022). The measure uses a score from 0 to 1, where 1 means the assistant wrote the entire text (reflecting complete reliance on the assistant) and 0 means the human wrote the entire text (no reliance). We use the term AI reliance for this measure, instead of mutuality, as we believe this concept better captures the construct used.
Inclusion. We draw on the concept of self-extension, which explains how objects become integral to one’s self-concept (Belk, 2013), to develop the notion of inclusion in AI-mediated communication. We focus on identity self-extension where the technology is a reflection of the user (Ross and Bayer, 2021; Mieczkowski, 2022). This literature suggests that if a user feels the technology reflects themselves, they will feel included. We therefore measure inclusion with the following statements: the writing assistant was made for people like me and the writing assistant sounded like I would write myself.
Control. Prior work has shown that feeling in control is not limited to one action or outcome. Feeling in control can encompass engaging in an action, control over the outcome, or both (Singh et al., 2022; Mieczkowski, 2022). For example, some users may a feel lack of control since they cannot choose which suggestions they are seeing (control during the writing process). However, the fact that they can select which suggestions are incorporated into the message can impact their sense of control over the process and the final version (Bhat et al., 2023). We therefore measure by asking participants how much control did you feel over the process of writing the message and how much control did you feel over the final version of the message (Mieczkowski, 2022).
Ownership. One aspect of ownership involves ”mineness”– the idea that a specific target belongs to an individual. The target can include physical objects but can also encompass intangible objects such as words and thoughts. In our experiment, we consider two targets– the message itself and the message style. Therefore, we also investigate ownership over the writing content and style. Our questions included to what extent do you feel like the message you wrote is yours and thinking back on the message writing activity, how much did the message sound like you (Mieczkowski, 2022).
4. Results
Our results expose the impact of writing style on perceptions of inclusion, control, and ownership and agency and hint at potential differences between genders in that respect. Our primary, preregistered analysis finds some significant effects of AI writing style on perceptions of control and ownership, but not on inclusion. A follow-up, exploratory analysis suggests that as AI reliance increases, perceptions of inclusion mediate the loss of agency. Additionally, the analysis suggests that heightened perceptions of inclusion lead to increased agency, particularly among minoritized genders.
We first provide an overview of the effects of using the AI auto-complete suggestions on the writing task. We examined three components of the writing task: the length of text written by participants, their writing time, and their degree of AI reliance. We performed a two-way ANOVA to compare the means between the three treatment groups (self-assured AI writing style, hesitant AI writing style, and control) and two gender groups (men and WNBSD participants). Overall, participants who wrote with the assistant wrote longer emails (M=160.18, SD=82.16) and spent less time on the task (M=303.82, SD=218.53) than participants in the control group (email length M=95.53, SD=35.64; time on task M=363.66, SD=292.68). The differences are statistically significant (length: ANOVA F=67.86, p¡0.00001; time: ANOVA F=6.35, p¡0.002). Post-hoc tests did not observe any differences between the two AI conditions in email length and time on task. In contrast, we did observe significant impacts of AI writing style on AI reliance: participants who wrote with the self-assured writing style model relied on the model more (M=0.55, SD=0.34) than participants who wrote with the hesitant writing style model (M=0.45, SD=0.35). These differences were also significant (t=3.16, p¡0.002). We did not observe gender effects or interaction effects between gender and the AI treatment on any of these metrics.
4.1. Inclusion
We examine the impact of AI writing style on our key measures, beginning with perceptions of inclusion. Figure 2 shows, for each gender condition (different bars) the frequency of different survey responses for the inclusion statements: the writing assistant was made for people like me (“Assistant Was Made for You” on the left), and the writing assistant sounded like I would write myself (“Assistant Sounds Like You” on the right). Figure 2a (top-left) shows that in the condition with the self-assured AI writing style, 68.3% of the men and 68.5% of the WNBSD participants either agreed or strongly agreed (the blue and dark blue shades on the right of the bars) that the writing assistant was made for people like them. As the figure also shows, in the hesitant AI writing style group, (the bottom-left Figure 2c), 65.1% of men and 62.9% of WNBSD participants either agreed or strongly agreed with this statement.
The responses to the other statement, whether the assistant aligned with the participants’ writing style, are shown on the right side of Figure 2. In general, participants more readily agreed with this statement compared to the earlier “made for you” statement. Figure 2b (top-right) shows the participants who wrote with the self-assured writing style model; of the participants who received this treatment, 50% of the men and 51.7% of the WNBSD participants either agreed or strongly agreed that the assistant sounded like how they would write themselves. In the condition with the hesitant AI writing style, (bottom-right Figure 2d), 42.9% of the men and 45.7% of the WNBSD participants either agreed or strongly agreed. A two-way ANOVA analysis revealed no significant effects for gender, AI treatment group, or the interaction between gender and group across all inclusion questions.
4.2. Control
Figure 3 illustrates, for each gender condition (different rows) the frequency of different survey responses for the following questions: how much control did you feel over the process of writing the message (“Control During Process” on the left), and how much control did you feel over the final version of the message (“Control Over Message” on the right). Figure 3a (top-left) shows in the condition with the self-assured AI writing style suggestions, during the writing process, 48% of men and 49% of WNBSD participants reported feeling a lot or great deal of control (indicated by the darker blue and darkest blue bars). In the condition with the hesitant AI writing style, seen in Figure 3c (bottom-left), 61.1% of the men and 61.2% of the WNBSD participants expressed a lot or a great deal of control during the writing process. A two-way ANOVA revealed that the difference between the AI treatment groups was significant (p¡0.05), meaning that participants in the hesitant condition reported more control over the process. There were no significant differences for gender, or interaction between gender and AI treatment group for this question.
In both treatment conditions, participants generally reported higher levels of control over the final version of the message than the writing process. In the condition with the self-assured AI writing style, (Figure 3b, top-right), 59.6% of men and 70.6% of WNBSD participants reported a lot or a great deal of control during the writing process. Figure 3d (bottom-right) shows 73% of men and 80.2% of WNBSD participants who wrote with a hesitant writing style model reported a lot or a great deal of control over the final version of the message. For this question as well, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p¡0.05), with participants in the hesitant AI writing style condition reported greater control over the final version of the message. At the same time, there were no significant differences in relation to gender or interactions between gender and AI treatment group for this question.
4.3. Ownership
Figure 4 illustrates the frequency of different survey responses for each gender condition (rows) for the following questions: to what extent do you feel like the message you wrote is yours (“Message Is Yours”), and thinking back on the message writing activity, how much did the message sound like you (“Message Sounds Like You”). In the self-assured AI writing style condition, (Figure 4a, top-left), 41.4% of men and 44.1% of WNBSD participants expressed a lot or a great deal of ownership over the message (as seen by the dark blue and darkest blue bars). In the hesitant AI writing style condition (Figure 4c, bottom-left), participants experienced a higher level of ownership with 55.6% of the men and 53.5% of the WNBSD participants expressing a similar degree of ownership over the message. A two-way ANOVA revealed that this difference between the AI treatment groups was significant (p¡0.05). There was no significant difference for gender, or interaction between gender and AI treatment group for this question.
Reflecting on the style of the message, in the self-assured AI writing style condition figures 4b and d on the right show similar trends: participants in the hesitant AI writing style condition (bottom) were more likely to say the final message reflected their style. Once again, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p¡0.05). There was no significant difference for gender, or interaction between gender and AI treatment group and gender for this question.
Overall, our results show AI writing style impacts control and ownership. Participants who wrote with the hesitant model, expressed greater control over the writing process and the final message and felt the message sounded more like them than participants who wrote with the self-assured model. However, our analysis did not reveal an impact of the AI writing style on perceived inclusion. Additionally, we did not see any effects of gender or interaction between gender and treatment group across inclusion, control, and ownership questions. We further explore these factors in the analysis below.
4.4. Exploratory Analysis: Inclusion and Agency
To better understand the relationship between the different measures and the participants’ interaction with the AI model, we perform an exploratory analysis that brings together multiple factors using a linear regression model. To perform this analysis, we combined various measures into simplified high-level constructs. First, we use as our dependent variable a compound measure of agency. The concept of agency has been defined in several different ways, but broadly captures the idea of “the capacity to alter a situation” (Neff and Nagy, 2018; Cesafsky et al., 2019). The variables that reflect that in our study were control and ownership. While distinct, the literature shows these concepts are closely related: whether a person has control greatly influences their perceived ownership (Mieczkowski, 2022; Braun et al., 2018). An analysis of our measures confirmed the control and ownership questions were indeed highly correlated in our data. We thus created a score for agency by performing a row-wise average across the control and ownership questions.
The independent variables in the model consisted of a simplified inclusion variable, the AI reliance score, the AI writing style, and demographic variables. To simplify the model and analysis, we combined the two highly correlated inclusion measures into one variable by performing a row-wise average across the two inclusion questions. The model also included the measure of AI Reliance as defined above, namely how much the participants accepted the AI suggestions in the writing process. Finally, we include in our model the AI writing style of the model that participants wrote with (the AI treatment), as well as more detailed demographic variables: age, gender, race, and income. The race and gender features were operationalized as binary variables where 1 reflected the privileged group and 0 reflected the minoritized group (e.g., for gender, 1=men, 0=WNBSD). We also mapped the AI writing style as a binary variable (0=hesitant model, 1=self-assured model). We operationalized the age and income demographic features as ordinal variables. Note again that we consider this analysis exploratory, as it was not part of the preregistered analysis presented above that focused on the main effects.
| Dependent variable: | |
| Agency | |
| Age | 0.024 |
| Race | 0.013 |
| Income | 0.023 |
| Gender | 0.571 |
| ‘AI Reliance‘ | 3.451 |
| Inclusion | 0.136 |
| ‘AI Writing Style‘ | 0.244 |
| Gender:‘AI Reliance‘ | 0.251 |
| ‘AI Reliance‘:Inclusion | 0.666 |
| Gender:Inclusion | 0.204 |
| Gender:‘AI Writing Style‘ | 0.050 |
| Constant | 3.661 |
| Observations | 489 |
| R | 0.354 |
| Adjusted R | 0.339 |
| Residual Std. Error | 0.715 (df = 477) |
| F Statistic | 23.747 (df = 11; 477) |
| Note: | p0.05; p0.01; p0.001 |
Table 1 shows the parameter estimates of a linear regression model predicting perceived agency. We tested several versions of the model that accounted for additional interactions (like AI writing style and AI reliance), and we present the model with the best fit. The model shows a main effect of AI writing style (Table 1, line 7) on agency. The negative value of the parameter signifies that writing with a self-assured model contributes to a loss of agency. We also see main effects of gender (line 4), showing men contributed to higher levels of agency. Perceptions of inclusion (line 6) also positively impacted agency, and AI reliance (line 5) had a negative impact. These effects were qualified by interaction effects between AI reliance and perceived inclusion (line 9) and gender and perceived inclusion (line 10), which we explore next.
We explore the interaction effects to better understand their impact on perceived agency. Figure 4(a) depicts the interaction between AI reliance and perceived inclusion. AI reliance is the x-axis and perceived agency is the y-axis. The graph compares the groups that showed high (dark blue), medium (blue), and low (light blue) levels of inclusion. Participants with a high level of inclusion, as indicated by the dark blue line in the figure, maintain their sense of agency even with increasing reliance on the writing assistant. Participants with a low (and to a lesser degree, medium) sense of inclusion lose their sense of agency rapidly when they rely more on AI.
Figure 4(b) explores the interaction between inclusion (x-axis) and gender. The graph compares men (blue line) and minoritized genders (WNBSD, light blue line). While inclusion and agency have a direct relationship for all gender identities, the slope for WNBSD participants is greater than the slope for men, indicating that perceived inclusion is more strongly tied to perceived agency for minoritized genders.
5. Discussion
Our findings enhance the current models of “writing with AI” by introducing new factors and furthering our understanding of their relationships. The findings also hint at the role that gender may play in such context, though this line of investigation requires additional substantiation. Overall, the findings offer implications for research and design of co-writing interactions with AI.
By introducing the construct of inclusion, our work extends the previous research on AI-MC which has examined perceptions of agency while interacting with an AI writing assistant (Lee et al., 2022; Mieczkowski, 2022; Mieczkowski et al., 2021) We measure the concept of inclusion after the co-writing process, by asking participants if they could envision themselves in the model and whether the model’s writing style matched their own. We show that this concept contributes to the perception of agency, a combined metric reflecting ownership and control over the writing process and the final product.
Informed by these findings, Figure 6 presents a new conceptual model of the factors contributing to agency in AI co-writing environments. This model is tentative: not all of these concepts were measured, let alone tested, in our work here. The model suggests that there are two paths for feeling agency in the co-writing process and its outcome. On the top, AI writing style affects AI-user alignment—which we define as the similarity between the AI writing style and the user’s own preferences, that may be derived from their identity. High AI-user alignment leads to increased perceived inclusion that will positively contribute to perceived agency. In our research, we did not measure AI-user alignment. Instead, we built on the literature suggesting that minoritized genders are less likely to be aligned with assertive AI suggestions – which was not the case in this study. At the bottom of the model, AI writing style affects AI-task alignment, which we define as the perceived usefulness of the AI assistant for the task (e.g. writing to ask for a promotion with an assertive AI). If there is high AI-task alignment, the user will rely on the model to complete the task. However, such reliance, as we have shown, can negatively contribute to the user’s perceived agency. The conceptual model includes an interaction between AI reliance and inclusion, as suggested by our exploratory results.
Our model raises several open questions that have implications for designers. The most salient question is the potential competition between AI-User alignment and AI-Task alignment. Or, how can developers strike a balance between models that are inclusive of various users but are also effective in completing the task? We note that even if the models align with both user and task, there is some conflict: the AI-User alignment will ultimately contribute to agency, but the AI alignment with the task, which will increase reliance, will reduce agency. This finding complicates the question of how developers can build models that help the user accomplish the task while also supporting the user’s agency, when there is tension between those.
The idea of inclusion via AI-User alignment captures our hypothesis that gender identity will play a role in perceived inclusion and agency. We hypothesized that minoritized genders would feel more included in the hesitant AI writing style condition, at least compared to men. However, our main findings did not show such gender effects. Although existing literature highlights language use differs by gender in negotiation contexts, it often relies on experimental methods (such as laboratory studies) that encompass factors beyond the scope of our experiment like job status, vocal pitch, or physical presence (Amanatullah and Morris, 2010; Mazei et al., 2015). Many of the laboratory studies show it is the small details, like the tone of voice or the wording, which makes gender differences salient (Thimm et al., 2008). The absence of physical cues, as shown in computer-mediated negotiations, reduces gender-based communication bias (Mazei et al., 2015).
While gender differences did not account for our expected main effect, our exploratory analysis also suggests there are still, potentially, some differences between genders. A noteworthy finding in this analysis was the significant interaction between gender and inclusion in relation to perceived agency. The interaction suggests that for minoritized genders, perceptions of inclusion have a greater effect on the perception of agency compared to men. If they hold, these findings lend support to our initial hypothesis: that people of different genders may react differently to writing assistants of different styles. This finding suggests that it may be useful to devise, develop, and test, language models that also optimize on generating perceptions of inclusion by people of different demographic backgrounds.